# SuperBPE: Space Travel for Language Models

### v2

@@ Abstract
**footnotetext: Equal contribution.The assumption across nearly all language model (LM) tokenization schemes is that tokens should be subwords, i.e., contained within word boundaries. While providing a seemingly reasonable inductive bias, is this common practice limiting the potential of modern LMs? Whitespace is not a reliable delimiter of meaning, as evidenced by multi-word expressions (e.g., by the way), crosslingual variation in the number of words needed to express a concept (e.g., spacesuit helmet in German is Raumanzughelm), and languages that do not use whitespace at all (e.g., Chinese). To explore the potential of tokenization beyond subwords, we introduce a “superword” tokenizer, SuperBPE, which incorporates a simple pretokenization curriculum into the byte-pair encoding (BPE) algorithm to first learn subwords, then superwords that bridge whitespace. This brings dramatic improvements in encoding efficiency: when fixing the vocabulary size to 200k, SuperBPE encodes a fixed piece of text with up to 33% fewer tokens than BPE on average. In experiments, we pretrain 8B transformer LMs from scratch while fixing the model size, vocabulary size, and train compute, varying only the algorithm for learning the vocabulary. Our model trained with SuperBPE achieves an average +4.0% absolute improvement over the BPE baseline across 30 downstream tasks (including +8.2% on MMLU), while simultaneously requiring 27% less compute at inference time. In analysis, we find that SuperBPE results in segmentations of text that are more uniform in per-token difficulty. Qualitatively, this may be because SuperBPE tokens often capture common multi-word expressions that function semantically as a single unit. SuperBPE is a straightforward, local modification to tokenization that improves both encoding efficiency and downstream performance, yielding better language models overall.111Code and artifacts are available at https://superbpe.github.io/.
@@ 1 Introduction
Tokenizers are the lens through which language models (LMs) view data: they segment a stream of bytes into a sequence of tokens in the LM vocabulary. In the era of transformer LMs, tokenization is done at the level of subwords, meaning that tokens consist of parts of words (including complete words), but cannot bridge whitespace. Intuitively, subword tokens capture meaningful and composable semantic units.
While seemingly reasonable, is this common practice a good one? Whitespace is an unreliable delimiter of meaning (Martin, 2017)—many groups of words (e.g., a lot of or search engine) function semantically as single units, and English speakers store thousands of such multi-word expressions in their mental lexicon (Church, 2011; Contreras Kallens & Christiansen, 2022). Crosslingually, there is considerable variation in whether a given meaning is conveyed by a single word or several words. At the extreme, languages such as Chinese and Japanese do not use whitespace at all, so that tokens in these languages can span multiple words or even entire sentences (e.g., in the tokenizers of Gpt-4o [OpenAI, 2024] or DeepSeekv3 [DeepSeek-AI et al., 2025]), yet this has seemingly not hindered LMs from performing well on these languages. Including multi-word tokens promises to be beneficial in several ways: it can lead to shorter token sequences, lowering the computational costs of LM training and inference, and may also offer representational advantages by segmenting text into more semantically cohesive units (Salehi et al., 2015; Otani et al., 2020; Hofmann et al., 2021).
In this work, we introduce a superword tokenization algorithm that produces a vocabulary of both subword and “superword” tokens, which we use to refer to tokens that bridge more than one word. Our method, SuperBPE, introduces a pretokenization curriculum to the popular byte-pair encoding (BPE) algorithm (Sennrich et al., 2016): whitespace pretokenization is initially used to enforce learning of subword tokens only (as done in conventional BPE), but is disabled in a second stage, where the tokenizer transitions to learning superword tokens. Notably, SuperBPE tokenizers scale much better with vocabulary size—while BPE quickly hits a point of diminishing returns and begins adding increasingly rare subwords to the vocabulary, SuperBPE can continue to discover common word sequences to treat as single tokens and improve encoding efficiency (see Figure 1).
In our main experiments, we pretrain English LMs at 8B scale from scratch. When fixing the model size, vocabulary size, and training compute—varying only the algorithm for learning the vocabulary—we find that models trained with SuperBPE tokenizers consistently and significantly improve over counterparts trained with a BPE tokenizer, while also being 27–33% more efficient at inference time. Our best SuperBPE model achieves an average +4.0% improvement over 30 downstream tasks, including +8.2% on MMLU, and wins on 25 of the 30 individual tasks (Table 1).
@@ 2 SuperBPE
We first explain the standard byte-pair encoding (BPE; Sennrich et al., 2016) tokenization algorithm (§2.1), then introduce SuperBPE, which extends BPE to superword tokens (§2.2).
@@ 2.1 Background on BPE
BPE is a tokenization algorithm that greedily learns a subword vocabulary given training data.222The algorithm originated in 1994 in the field of data compression (Gage, 1994). The algorithm takes a sample of text and a target vocabulary size as input. Note that although the creation of a tokenizer is referred to as “learning,” there are no parameters involved in the case of BPE, and the algorithm is completely deterministic given the data.
The first step of BPE is pretokenization, which splits the text into chunks that limit the extent of tokenization: merges cannot bridge these chunks, so the final learned tokens will be parts of these chunks. Canonically, pretokenization in BPE consists of splitting on whitespace, so that common word sequences do not become a single token. This made sense in the historical context of Sennrich et al. (2016), which aimed to improve word-level tokenization by segmenting words into morphologically meaningful subwords.
After pretokenization, the iterative learning algorithm begins. The training text is first split into bytes; the starting vocabulary is the set of all bytes. Then, the frequencies of all pairs of neighboring tokens are recorded, and the most frequent pair is merged into a single, new token at every position in the text where it occurs. The newly merged token is added to the vocabulary. For instance, if the merge is (t, he), then all instances of the token sequence [t, he] will be replaced with the, which is added to the vocabulary. The token pair frequencies are then updated, and the next most frequent pair is again merged into a new token. This continues until vocabulary reaches the target size .
@@ 2.2 SuperBPE tokenization
SuperBPE introduces a simple intervention on the pretokenization step, separating tokenizer training into two discrete phases so that the tokenizer first learns subwords (by using pretokenization to prevent merges across whitespace) and then superwords (by lifting this restriction). Stage 1 is equivalent to regular BPE training and continues up to a certain vocabulary size , which we call the transition point (). In stage 2, tokenizer training resumes from the vocabulary learned so far, but this time whitespace pretokenization is skipped. As a result, token pairs bridging whitespace are considered, making it possible for superwords to be added to the vocabulary. Intuitively, we wish for our tokenizer to first learn base units of semantic meaning, then combine these units into common sequences for a much more efficient vocabulary. Note that corresponds to BPE, and corresponds to a naive revision of BPE that foregoes whitespace pretokenization at any point in training.
@@
We note that training tokenizers requires more system memory and CPU time when done without whitespace pretokenization (as in stage 2 of SuperBPE). This is because the training data is typically represented by a dictionary of “words” along with their counts. With whitespace pretokenization, the “words” are whitespace-separated chunks (e.g., common words) stored once alongside a large count, granting substantial savings in memory. Without whitespace pretokenization, the “words” are extremely long (e.g., entire training documents), leading to minimal deduplication of the text and excessively large dictionaries. Fortunately, tokenizer training must be done only once; in our experiments, SuperBPE tokenizers train in a few hours on 100 CPUs, a negligible cost compared to LLM pretraining.
@@ 2.3 Encoding efficiency
A tokenizer’s encoding efficiency can be measured in bytes-per-token, i.e., how many UTF-8 bytes are encoded, on average, in each token over a large corpus of text (see calculation in Figure 1). We train a series of tokenizers on a 10 GB subset of data from Olmo 2’s pretraining corpus, and evaluate encoding efficiency on a held-out subset.
Shown in Figure 1, SuperBPE scales much better with vocabulary size than BPE. BPE quickly plateaus around a vocabulary size of 50K, achieving 4.45 bytes-per-token at a vocabulary size of 200k. In fact, even with infinite vocabulary size (namely, if every whitespace-delimited word were in the vocabulary), BPE cannot exceed 4.68 bytes-per-token, i.e., the average word length in the held-out subset. SuperBPE exceeds this upper bound with a mere 12k vocabulary size, and reaches 5.55 bytes-per-token at 50K and 6.63 at 200k.
Perhaps surprisingly, SuperBPE is also more efficient than BPE with whitespace pretokenization completely disabled. Since BPE is a greedy algorithm, completely disabling whitespace pretokenization may cause BPE to make highly suboptimal choices early on. In particular, tokens in this setting often consist of the end of the previous word and start of the next word, as opposed to sequences of complete words. By keeping whitespace pretokenization on at the beginning, we can avoid suboptimal choices while still obtaining a tokenizer with superwords.
@@ 3.1 Setup
We first pretrain 8B models with BPE and SuperBPE tokenizers. We use the Olmo2 7B (OLMo et al., 2024) training configuration,333Olmo2 7B has 7.30B parameters while our 8B BPE and SuperBPE models have 8.12B parameters due to their increased vocabulary size. including the model architecture, training hyperparameters, and pretraining corpus, but reduce the total number of training steps to correspond to 330B tokens (compared to 4T) for the sake of compute. Following prior work (Pagnoni et al., 2024), we also fix the effective context size (measured in bytes) for each model. This prevents SuperBPE models from gaining an advantage by seeing more textual context for the same next-token prediction (Xiong et al., 2024). As more efficient models will have a shorter context length in tokens, the training steps are adjusted accordingly to match the total train FLOPs at the end of training.444In practice, models using our more efficient tokenizers could shift some or all of the “saved” context FLOPs to longer effective contexts instead of more training steps. Note that in this setting, a same-sized SuperBPE model uses fewer inference FLOPs compared to the BPE model.
We fix the vocabulary size of all tokenizers to 200,000 (in the same ballpark as e.g., Gemma at 250k [Google, 2024], Gpt-4o at 200k, Llama3 at 130k [Meta, 2024]).555For 8B models, 200k vocabulary size is close to the recommendation of Tao et al. (2024) based on primarily-English data. We fix the vocabulary size to simplify comparisons between models. We consider three transition points for SuperBPE: 80k, which has the best encoding efficiency, and two later transitions, 160k and 180k. All tokenizers are trained on the same 10 GB subset of Olmo2’s pretraining mix. We provide further details about tokenizer training in §A.1.
We also train a slightly larger 11B SuperBPE model with 180k which approximately matches the 8B BPE baseline in total bytes of training data seen as well as both train and inference compute. See Table 2 for exact specifications for all runs.
@@ 3.2 Results on downstream tasks
We evaluate on 30 benchmarks covering knowledge, math & reasoning, coding, reading comprehension, commonsense, language understanding, and string manipulation. The full evaluation suite is shown in Table 1; evaluation details can be found in §A.3.
Figure 3 shows the task average during pretraining. All SuperBPE models substantially outperform the BPE baseline at the end of training. The strongest 8B SuperBPE model, which has transition point 180k (the latest one we consider), outperforms the baseline by 4.0% on average, and wins on 25 out of 30 individual tasks. Table 1 shows the per-task performance for this model (see §A.3 for results for the other models). The largest gains appear to be on multiple choice tasks; when considering these alone, the performance improvement is +9.7%. The only task on which SuperBPE loses in a statistically significant way is LAMBADA; here, we observe that SuperBPE is actually ahead for the majority of training checkpoints, but accuracy dips at the end from 75.8% to 70.6% (see Figure 11).
Note that while the choice of transition point affects the performance of the resulting model, all reasonable choices are significantly stronger than the baseline. When using the most encoding-efficient transition point of 80k, we see a +3.1% task improvement over BPE while reducing inference compute by 35%. Later transition points empirically give up some gains in encoding efficiency in exchange for further improved performance.666This finding adds to the ongoing debate about the relationship between tokenization compression and LM performance (Gallé, 2019; Goldman et al., 2024; Schmidt et al., 2024), providing further evidence that a higher compression does not necessarily lead to better performance.
@@ 4.1 Language modeling
Following prior work (Biderman et al., 2023; Xue et al., 2022; Yu et al., 2023; Wang et al., 2024), we evaluate language modeling performance using bits-per-byte (BPB), which normalizes the loss by the encoding efficiency of the tokenizer to fairly compare models with different tokenizers. This is necessary because longer tokens, on average, contain more information and therefore are more difficult to predict. Bits-per-byte is defined as where is the length of the text in bytes and is the sum of the cross-entropy loss over the entire text.777Bits-per-byte of different models are considered comparable because total cross-entropy loss is a universal quantity representing the number of additional bits required to reconstruct the text given the model. This quantity is normalized by the number of bytes for easier interpretation. We find that BPE 8B, SuperBPE 8B ( 180k), and SuperBPE 11B attain 0.7465, 0.7482, and 0.7445 BPB, respectively, at the end of training. Though the numbers are all close, the rankings of the models according to BPB and downstream task performance are not consistent.
@@
4.2 Fine-grained analysis of bits-per-byte
Why does the SuperBPE 8B model achieve slightly higher normalized language modeling loss (§4.1) than the baseline BPE model, despite outperforming it on a wide variety of downstream tasks (§3.2)? To investigate this, we plot the distribution of per-token BPB888The per-token BPB is the per-token loss (in bits) divided by the average encoding efficiency. for both models on data sampled from the pretraining data mixture in Figure 4.
We see that, although the BPE and SuperBPE models have very similar BPB on average, that loss is distributed very differently over the training data. Compared to the baseline, the SuperBPE model makes fewer predictions with either very high or very low loss.

### v3

@@ Abstract
**footnotetext: Equal contribution.The assumption across nearly all language model (LM) tokenization schemes is that tokens should be subwords, i.e., contained within word boundaries. Despite providing a seemingly reasonable inductive bias, we question whether this common practice limits the potential of modern LMs. Whitespace is not a reliable delimiter of meaning, as evidenced by multi-word expressions (e.g., by the way), cross-lingual variation in the number of words needed to express a concept (e.g., spacesuit helmet in German is raumanzughelm), and languages that do not use whitespace at all (e.g., Chinese). To explore the potential of tokenization beyond subwords, we introduce a “superword” tokenizer, SuperBPE, that incorporates a simple pretokenization curriculum into the byte-pair encoding (BPE) algorithm to first learn subwords and then superwords that bridge whitespace. This modification dramatically improves encoding efficiency: when limiting vocabulary size to 200k, SuperBPE encodes a fixed piece of text with up to 33% fewer tokens on average than BPE. In experiments, we pretrain 8B transformer LMs from scratch while fixing model size, vocabulary size, and train compute, varying only the algorithm for learning the vocabulary. Our model trained with SuperBPE achieves an average +4.0% absolute improvement over the BPE baseline across 30 downstream tasks (including +8.2% on MMLU), while simultaneously requiring 27% less compute at inference time. In analysis, we find that SuperBPE produces segmentations of text that are more uniform in per-token difficulty, perhaps because SuperBPE tokens often capture common multi-word expressions that function semantically as a single unit. In sum, SuperBPE offers a straightforward and local modification to tokenization that improves both encoding efficiency and downstream performance, yielding better LMs overall.111Code and artifacts are available at https://superbpe.github.io/.
@@ 1 Introduction
Tokenizers are the lens through which language models (LMs) view data: they segment a stream of bytes into a sequence of tokens in the LM vocabulary. In the era of transformer LMs, tokenization is done at the level of subwords, meaning that tokens consist of parts of words (including complete words), but they cannot bridge whitespace. Intuitively, subword tokens capture meaningful and composable semantic units.
Although seemingly reasonable, is this common practice a good one? Whitespace is an unreliable delimiter of meaning (Martin, 2017); many groups of words (e.g., a lot of or search engine) function semantically as single units, and English speakers store thousands of such multi-word expressions in their mental lexicon (Church, 2011; Contreras Kallens & Christiansen, 2022). Cross-lingually, there is considerable variation in whether a given meaning is conveyed by a single word or several words. At the extreme, languages such as Chinese and Japanese do not use whitespace at all, and tokens in these languages can span multiple words or even entire sentences (e.g., the tokenizers of Gpt-4o [OpenAI, 2024] or DeepSeekv3 [DeepSeek-AI, 2025]), but this has seemingly not hindered LMs from performing well on these languages. In fact, including multi-word tokens promises to be beneficial in many ways: it may shorten token sequences, lowering the costs of LM training and inference, and offer representational advantages by segmenting text into more semantically cohesive units (Salehi et al., 2015; Otani et al., 2020; Hofmann et al., 2021).
In this work, we introduce a superword tokenization algorithm that produces a vocabulary of both subword and “superword” tokens, which we use to describe tokens bridging more than one word. Our method, SuperBPE, introduces a pretokenization curriculum to the popular byte-pair encoding (BPE) algorithm (Sennrich et al., 2016): whitespace pretokenization is initially used to enforce learning of subword tokens only (as done in conventional BPE), but it is disabled in a second stage, where the tokenizer transitions to learning superword tokens. Notably, SuperBPE tokenizers scale much better with vocabulary size: BPE quickly hits a point of diminishing returns and begins adding increasingly rare subwords to the vocabulary, while SuperBPE continues to discover common word sequences to treat as single tokens and improve encoding efficiency (see Figure 1).
In our experiments, we pretrain English LMs at 8B scale from scratch. When fixing the model size, vocabulary size, and training compute—varying only the algorithm for learning the vocabulary—we find that models trained with SuperBPE tokenizers consistently and significantly improve over counterparts trained with a BPE tokenizer while also being 27% to 33% more efficient at inference time. Our best SuperBPE model achieves an average improvement of +4.0% over 30 downstream tasks, including +8.2% on MMLU, and wins on 25 of the 30 individual tasks (Table 1).
@@ 2 SuperBPE
We first explain the standard byte-pair encoding (BPE; Sennrich et al., 2016) tokenization algorithm (§2.1), and then introduce SuperBPE, which extends BPE to superwords (§2.2).
@@ 2.1 Background on BPE
BPE is a tokenization algorithm that greedily learns a subword vocabulary given training data.222The algorithm originated in 1994 in the field of data compression (Gage, 1994). The algorithm takes a sample of text and a target vocabulary size as input.333Note that although the creation of a tokenizer is referred to as “learning,” there are no parameters involved in the case of BPE, and the algorithm is completely deterministic given the data.
The first step of BPE is pretokenization, which splits the text into chunks that limit the extent of tokenization; merges cannot bridge these chunks, so the final learned tokens are parts of these chunks. Canonically, pretokenization in BPE consists of splitting on whitespace so that common word sequences do not become a single token. This made sense given the historical context of Sennrich et al. (2016), which aimed to improve word-level tokenization by segmenting words into morphologically meaningful subwords.
After pretokenization, the iterative learning algorithm begins. Training text is first split into bytes; the starting vocabulary is the set of all bytes. Then, the frequencies of all pairs of neighboring tokens are recorded, and the most frequent pair is merged into a single, new token at every position in the text where it occurs. The newly merged token is added to the vocabulary. For instance, if the merge is (t, he), then all instances of the token sequence [t, he] will be replaced with the, which is added to the vocabulary. The token pair frequencies are then updated, and the next most frequent pair is again merged into a new token. This continues until the vocabulary reaches the target size .
@@ 2.2 SuperBPE tokenization
SuperBPE introduces a simple intervention in the pretokenization step, separating tokenizer training into two discrete phases, wherein the tokenizer (1) first learns subwords (by using pretokenization to prevent merges across whitespace) and then (2) learns superwords (by lifting this restriction). Stage 1 is equivalent to regular BPE training and continues up to a certain vocabulary size , which we call the transition point (). In stage 2, tokenizer training resumes from the vocabulary learned thus far, but this time whitespace pretokenization is skipped. As a result, token pairs that bridge whitespace are considered, enabling superwords to be added to the vocabulary. Intuitively, we intend for our tokenizer to first learn base units of semantic meaning, then combine these units into common sequences for a much more efficient vocabulary. Note that corresponds to BPE, and corresponds to a naive revision of BPE that foregoes whitespace pretokenization at any point in training.
@@
We note that training tokenizers requires more system memory and CPU time when done without whitespace pretokenization (as in stage 2 of SuperBPE). This is because the training data is typically represented by a dictionary of “words” along with their counts. With whitespace pretokenization, the “words” are whitespace-separated chunks (e.g., common words) stored once along with a large count, conferring substantial savings in memory. Without whitespace pretokenization, the “words” are extremely long (e.g., entire training documents), leading to minimal deduplication of the text and excessively large dictionaries. Fortunately, tokenizer training must be done only once; in our experiments, SuperBPE tokenizers train in a few hours on 100 CPUs, a negligible cost compared to LLM pretraining.
@@ 2.3 Encoding efficiency
A tokenizer’s encoding efficiency can be measured in bytes-per-token, i.e., how many UTF-8 bytes are encoded, on average, in each token over a large corpus of text (see calculation in Figure 1). We train a series of tokenizers on a 10 GB subset of data from Olmo 2’s pretraining corpus and evaluate encoding efficiency on a held-out subset.
Shown in Figure 1, SuperBPE scales much better with vocabulary size than does BPE. BPE quickly plateaus around a vocabulary size of 50K, achieving 4.45 bytes-per-token at a vocabulary size of 200k. In fact, even with infinite vocabulary size (namely, if every whitespace-delimited word were in the vocabulary), BPE cannot exceed 4.68 bytes-per-token, i.e., the average word length in the held-out subset. SuperBPE exceeds this upper bound with a mere 12k vocabulary size and reaches 5.55 bytes-per-token at 50K and 6.63 at 200k.
Surprisingly, SuperBPE is also more efficient than BPE with whitespace pretokenization completely disabled. Since BPE is a greedy algorithm, completely disabling whitespace pretokenization may cause it to make highly suboptimal choices early on. In particular, tokens in this setting often consist of the end of the previous word and start of the next word, as opposed to sequences of complete words. By keeping whitespace pretokenization on at the beginning, we can avoid suboptimal choices while still obtaining a tokenizer with superwords.
@@ 3.1 Setup
We first pretrain 8B models with BPE and SuperBPE tokenizers. We use the Olmo2 7B (OLMo et al., 2024) training configuration,444Olmo2 7B has 7.30B parameters, while our 8B BPE and SuperBPE models have 8.12B parameters due to their increased vocabulary size. including the model architecture, training hyperparameters, and pretraining corpus, but reduce the total number of training steps to correspond to 330B tokens (compared to 4T). Following prior work (Pagnoni et al., 2024), we also fix the effective context size (measured in bytes) for each model. This prevents SuperBPE models from gaining an advantage by seeing more textual context for the same next-token prediction (we provide analysis on this in §B.1). Since more efficient models have a shorter context length in tokens, the training steps are adjusted accordingly to match the total train FLOPs at the end of training.555In practice, models using our more efficient tokenizers could shift some or all of the “saved” context FLOPs to longer effective contexts instead of more training steps. Note that in this setting, a same-sized SuperBPE model uses fewer inference FLOPs than the BPE model.
We fix the vocabulary size of all tokenizers to 200,000 (in the same ballpark as, e.g., Gemma at 250k [Google, 2024], Gpt-4o at 200k, and Llama3 at 130k [Meta, 2024]).666For 8B models, a 200k vocabulary size is close to the recommendation of Tao et al. (2024) based on primarily English data. We fix the vocabulary size to simplify comparisons between models. We consider three transition points for SuperBPE: 80k, which has the best encoding efficiency, and two later transitions, 160k and 180k. All tokenizers are trained on the same 10 GB subset of Olmo2’s pretraining mix. §A.1 provides further details about tokenizer training.
We also train a slightly larger 11B SuperBPE model with 180k, which approximately matches the 8B BPE baseline in total bytes of training data seen as well as both train and inference compute. See Table 2 for exact specifications for all runs.
@@ 3.2 Results on downstream tasks
We evaluate SuperBPE on 30 benchmarks covering knowledge, math & reasoning, coding, reading comprehension, common sense, language understanding, and string manipulation. The full evaluation suite is shown in Table 1 and evaluation details are in §A.4.
Figure 3 shows the task average during pretraining. All SuperBPE models substantially outperform the BPE baseline at the end of training. The strongest 8B SuperBPE model, which has transition point 180k (the latest one we consider), outperforms the baseline by 4.0% on average and wins on 25 of 30 individual tasks. Table 1 shows the per-task performance for this model (see §A.4 for results for the other models). The largest gains are on multiple choice tasks; when considering these alone, the performance improvement grows to +9.7%. The only task on which SuperBPE loses in a statistically significant way is LAMBADA; here, we observe that SuperBPE is actually ahead for the majority of training checkpoints, but accuracy dips at the end from 75.8% to 70.6% (see Figure 12).
Notably, while the choice of transition point affects the performance of the resulting model, all reasonable choices are significantly stronger than the baseline. When using the most encoding-efficient transition point, i.e., 80k, we see a +3.1% task improvement over BPE and inference compute reduced by 35%. Later transition points empirically cede some gains in encoding efficiency in exchange for further improvements in performance.777This finding adds to the ongoing debate about the relationship between tokenization compression and LM performance (Gallé, 2019; Goldman et al., 2024; Schmidt et al., 2024), providing further evidence that higher compression does not necessarily improve performance.
@@ 4.1 Language modeling
Following prior work (Biderman et al., 2023; Xue et al., 2022; Yu et al., 2023; Wang et al., 2024), we evaluate language modeling performance using bits-per-byte (BPB), which normalizes the loss by the tokenizer’s encoding efficiency to fairly compare models with different tokenizers. This is necessary because longer tokens, on average, contain more information and therefore are more difficult to predict. Bits-per-byte is defined as where is the length of the text in bytes and is the sum of the cross-entropy loss over the entire text.888Bits-per-byte of different models are considered comparable because total cross-entropy loss is a universal quantity representing the number of additional bits required to reconstruct the text given the model. This quantity is normalized by the number of bytes for easier interpretation. We find that BPE 8B, SuperBPE 8B ( 180k), and SuperBPE 11B attain 0.7465, 0.7482, and 0.7445 BPB, respectively, at the end of training. Although these numbers do not differ appreciably, the ranking of models according to BPB and downstream task performance are not consistent.
@@
4.2 Loss distribution analysis
Why does the SuperBPE 8B model achieve slightly higher normalized language modeling loss (§4.1) than the baseline BPE model despite outperforming it on a wide variety of downstream tasks (§3.2)? To investigate this, we plot the distribution of per-token BPB999The per-token BPB is the per-token loss (in bits) divided by the average encoding efficiency. for both models on data sampled from the pretraining data mixture in Figure 4.
Although the BPE and SuperBPE models have very similar BPB on average, we see that loss is distributed very differently over the training data. Compared to the baseline, the SuperBPE model makes fewer predictions with either very high or very low loss.